OpenStudy (anonymous):

Need HELP PLEASE!! Create quadratic equation and demonstrate how it would be solved by graphing, factoring, the quadratic formula, and by completing the square. (5 points)

OpenStudy (anonymous):

I have no idea....Math is one of weaknesses. I'm sooo sorry

OpenStudy (anonymous):

^

OpenStudy (anonymous):

;(

OpenStudy (anonymous):

im sorry....google it???

OpenStudy (anonymous):

ok i found something :D

OpenStudy (anonymous):

ok :)

OpenStudy (nubeer):

hmm create a quadratic equation? lol means can take any?

OpenStudy (anonymous):

what shy said ;D

OpenStudy (anonymous):

x^2-6x=0

OpenStudy (anonymous):

by factoring:x(x-6)=0 x=0 or x-6=0,x=6 by completing the square: x^2-6x+(-6/2)^2=(-6/2)^2 x^2-6x+(-6/2)^2=9 (x-3)^2=9 x-3=3 ,x=6 or x-3=-3,x=0

OpenStudy (anonymous):

by the quadratic formula:

OpenStudy (anonymous):

\[x=\frac{ -6\pm \sqrt{36} }{ 2}=0 or 6\]

OpenStudy (anonymous):